OpenStudy (anonymous):

a plumber's bill b is based on $125 for materials and $50 per hour for t hours of labor. this situation can be represented by the function rule b=50t + 125. suppose the plumber works for 3 1/4h. how much is the bill?

OpenStudy (debbieg):

Just plug in 3 1/4 for t, then evaluate that expression to get b: \[b=50\left( 3\frac{ 1 }{ 4 } \right)+125\] You might find it easier to write the mixed number as a number with a decimal part.

OpenStudy (anonymous):

I do not understand @DebbieG

OpenStudy (debbieg):

Which part? Making the substitution in the formula? Or how to then evaluate that expression on the right hand side?

OpenStudy (debbieg):

The equation gives you b (amount of bill) for any amount of time t. So you plug in the amount of time for t in the equation, and then when you evaluate that, you will get the amount of the bill.

OpenStudy (anonymous):

but do not know how, do

OpenStudy (debbieg):

I'm.... not sure what that means. But I'll show you an example, using a different value for t. Then you can try to do your actual problem on your own. Let's suppose the plumber works for 2 and 3/4 hours, so \(\Large t=2\dfrac{3}{4}\). Now substituting that back into my equation: \(\Large b=50\left( 2\dfrac{ 3 }{ 4 } \right)+125\) Now since we are computing money, it's probably easiest to just convert that mixed number to a decimal. So since 3/4=0.75, we have \(\Large b=50\left( 2.75 \right)+125\) Now, depending on what you are allowed, you can either do that product by hand or use your calculator. Either way, you should get: \(\Large b=137.5+125\) and then just add those to get \(\Large b=262.5\) so the bill is \(\Large $262.50\) Now you try it, with your value for t.

OpenStudy (anonymous):

is $287.50

OpenStudy (debbieg):

Very good! You got it! :)